A pharmacist wishes to mix a solution that is 5​% Minoxidil. She has on hand 80 ml of a 2​% solution and wishes to add some 7​% solution to obtain the desired 5​% solution. How much 7​% solution should she​ add?
Answer by Theo(13342)

she wants to get a mix of 5% minoxidil.

she has 80 ml of 2% solution of minoxidil.

she has a certain quantity of 7% minoxidil solution on hand that she can use.

let x equal the amount of 7% solution that she has to add.

let y = the total amount of solution that she will wind up with.

you have two equations that need to be solved simultaneously.

they are:

total solution:

80 + x = y

total amount of minoxidil:

.02 * 80 + .07 * x = .05 * y

since y = 80 + x, then replace y with 80 + x in the second equation to get:

.02 * 80 + .07 * x = .05 * (80 + x)

simplify this equation to get:

1.6 + .07 * x = 4 + .05 * x

subtract 1.6 from both sides of the equation and subtract .05 * x from both sides of the equation to get:

.07 * x - .05 * x = 4 - 1.6

simplify to get .02 * x = 2.4

divide both sides of the equation by .02 to get x = 2.4 / .02 = 120.

she would need 120 ml of the 7% solution to get a 5% solution.

the total number of mls would be 80 + 120 = 200.

the total amount of minoxidil would be .02 * 80 + .07 * 120 = 1.6 + 8.4 = 10.

the percent of minoxidil in the final solution would be 10 / j200 = .05 * 100 = 5%.